OpenStudy (anonymous):

Find the value of x: 3x/4 = -24 x = 32 x = 1/32 x= -32 x = -1/32

OpenStudy (idealist10):

3x/4=-24 3x=-24*4 3x=-96 x=-96/3 x=-32 Answer: -32

OpenStudy (anonymous):

answer is -32